# hdMTD: An R Package for High-Dimensional Mixture Transition Distribution Models

### v1

@@ 1. Apply the CUT algorithm to the FS output:
[1] 364   6   1 In this example, the CUT method retained all three lags, indicating that it considers all elements in S = c(1, 6, 364) to be relevant.
@@ 2. Use BIC-based model selection on the FS output:
       1             1,364           1,6,364    smallest: 1,6,3641720.801          1690.543          1674.080             1674.080 This performs an exhaustive search over the $2^{3}-1=7$ possible non-empty subsets of S = c(1, 6, 364), selecting the configuration with the lowest Bayesian information criterion (BIC). Since byl = TRUE, the output reports the best-performing subset (top row) and its corresponding BIC value (bottom row) for each possible subset size (from 1 to 3 lags). In this example, the full set c(1, 6, 364) yields the lowest BIC, and is thus selected as the optimal lag configuration.
@@
R> P_FS <- probs(Temp12_Train, S = c(1, 6, 364), matrixform = T)R> P_FS
             1         2111 0.86626140 0.1337386112 0.24736842 0.7526316121 0.77157360 0.2284264122 0.13318777 0.8668122211 0.78846154 0.2115385212 0.10972569 0.8902743221 0.57506361 0.4249364222 0.07283555 0.9271645 Recall that the symbols in the rows represent occurrences at the pasts in $S$, ordered from oldest to most recent. For example sequence $112$ (in the second row) means that symbol $1$ occurred at lag 364-364, symbol $1$ at lag 6-6 and symbol $2$ at lag 1-1.
@@
R> P_MC2 <- probs(Temp12_Train, S = c(1, 2), matrixform = TRUE)R> P_MC2
            1         211 0.77813505 0.221865012 0.16326531 0.836734721 0.60233918 0.397660822 0.09064976 0.9093502 We now compare the predictive capabilities of this classical selection method with our proposed method (FS). Additionally, we include a naive baseline model (Ind), which assumes independence between observations: each symbol is drawn according to the overall empirical distribution of the training set.
@@
R and all associated packages are publicly available via the Comprehensive R Archive Network (CRAN) at https://CRAN.R-project.org/. All analyses in this study were performed using R version 4.4.1 (R Core Team, 2024) with the hdMTD package (v0.1.0), along with the following auxiliary packages: dplyr (Wickham et al., 2023a), ggplot2 (Wickham, 2016), lubridate (Grolemund and Wickham, 2011), purrr (Wickham et al., 2023b), and tidyr (Wickham et al., 2024).
@@ 6 Conclusions
Given a sample from an MTD model, hdMTD can retrieve the set of relevant pasts (lags) even when the order of the model is proportional to the sample size. Our implementation of the lag selection method can be seen as a feature selection method for non-parametric categorical time series and fills an important gap in the literature. The key feature of our lag selection method is that we can estimate the set of candidates for relevant lags a priori without the need to estimate the $d$-dimensional joint probability measure of a Markov chain of order $d$. Once the set of candidates is estimated, we apply an adaptive thresholding to eliminate the lags that were not relevant. Our method is provably consistent for a wide range of conditions (Ost and Takahashi, 2023).
The hdMTD package also implements a computationally efficient perfect (exact) sampling algorithm for MTD models. The main advantage of the perfect sampling algorithm is that it does not require choosing hyperparameters like burn-in duration, and it is not necessary to verify whether the obtained sample is stationary, as it is guaranteed to generate stationary samples. The exact sampling algorithm is especially relevant when the order of the MTD is large, as choosing hyperparameters or verifying the stationarity of the sample with long-range dependence can be hard. Perfect sampling algorithm for Markov chains of order $d$ is a theoretically well-studied topic, but a practical implementation seems to be lacking (Comets et al., 2002). To our knowledge, hdMTD is the first implementation of perfect simulation specific to MTD models. We expect that the toolset included in hdMTD will help with the analysis of stochastic phenomena with long-range dependencies that were previously impervious to rigorous statistical analysis.

### v2

@@ 1. Apply the CUT algorithm to the FS output:
[1] 1 6 364
In this example, the CUT method retained all three lags, indicating that it considers all elements in S = c(1, 6, 364) to be relevant.
@@ 2. Use BIC-based model selection on the FS output:
1 1,364 1,6,364 smallest: 1,6,3641720.801 1690.543 1674.080 1674.080
This performs an exhaustive search over the $2^{3}-1=7$ possible non-empty subsets of S = c(1, 6, 364), selecting the configuration with the lowest Bayesian information criterion (BIC). Since byl = TRUE, the output reports the best-performing subset (top row) and its corresponding BIC value (bottom row) for each possible subset size (from 1 to 3 lags). In this example, the full set c(1, 6, 364) yields the lowest BIC, and is thus selected as the optimal lag configuration.
@@
R> P_FS <- empirical_probs(Temp12_Train, S = c(1, 6, 364), matrixform = T)R> P_FS
1 2111 0.86626140 0.1337386112 0.24736842 0.7526316121 0.77157360 0.2284264122 0.13318777 0.8668122211 0.78846154 0.2115385212 0.10972569 0.8902743221 0.57506361 0.4249364222 0.07283555 0.9271645
Recall that the symbols in the rows represent occurrences at the pasts in $S$, ordered from oldest to most recent. For example sequence $112$ (in the second row) means that symbol $1$ occurred at lag 364-364, symbol $1$ at lag 6-6 and symbol $2$ at lag 1-1.
@@
R> P_MC2 <- empirical_probs(Temp12_Train, S = c(1, 2), matrixform = TRUE)R> P_MC2
1 211 0.77813505 0.221865012 0.16326531 0.836734721 0.60233918 0.397660822 0.09064976 0.9093502
We now compare the predictive capabilities of this classical selection method with our proposed method (FS). Additionally, we include a naive baseline model (Ind), which assumes independence between observations: each symbol is drawn according to the overall empirical distribution of the training set.
@@
R and all associated packages are publicly available via the Comprehensive R Archive Network (CRAN) at https://CRAN.R-project.org/. All analyses in this study were performed using R version 4.4.1 (R-base) with the hdMTD package (v0.1.3), along with the following auxiliary packages: dplyr (pkg:dplyr), future (future), future.apply (future), ggplot2 (pkg:ggplot2), lubridate (pkg:lubridate), purrr (pkg:purrr), and tidyr (pkg:tidyr).
@@ 6 Conclusions
Given a sample from an MTD model, hdMTD can retrieve the set of relevant pasts (lags) even when the order of the model is proportional to the sample size. Our implementation of the lag selection method can be seen as a feature selection method for non-parametric categorical time series and fills an important gap in the literature. The key feature of our lag selection method is that we can estimate the set of candidates for relevant lags a priori without the need to estimate the $d$-dimensional joint probability measure of a Markov chain of order $d$. Once the set of candidates is estimated, we apply an adaptive thresholding to eliminate the lags that were not relevant. Our method is provably consistent for a wide range of conditions (Ost&Takahashi).
The hdMTD package also implements a computationally efficient perfect (exact) sampling algorithm for MTD models. The main advantage of the perfect sampling algorithm is that it does not require choosing hyperparameters like burn-in duration, and it is not necessary to verify whether the obtained sample is stationary, as it is guaranteed to generate stationary samples. The exact sampling algorithm is especially relevant when the order of the MTD is large, as choosing hyperparameters or verifying the stationarity of the sample with long-range dependence can be hard. Perfect sampling algorithm for Markov chains of order $d$ is a theoretically well-studied topic, but a practical implementation seems to be lacking (Comets2002Proceses). To our knowledge, hdMTD is the first implementation of perfect simulation specific to MTD models. We expect that the toolset included in hdMTD will help with the analysis of stochastic phenomena with long-range dependencies that were previously impervious to rigorous statistical analysis.